OpenStudy (trojanpoem):

\[x^{x!} = x!^{x}\] find the values of x Using x = x! is refused. Any help on this ? ++ the legendary mathematical problem

OpenStudy (amistre64):

logs maybe?

ganeshie8 (ganeshie8):

1, 2 are trivial solutions for x>=3 show that x^x! > x!^x

ganeshie8 (ganeshie8):

yes, but induction might be tricky..

OpenStudy (loser66):

to me, it is good. I tried and got it

OpenStudy (trojanpoem):

The answer is 2 , 1 but I am looking for steps.

ganeshie8 (ganeshie8):

me too

OpenStudy (trojanpoem):

You are also looking for the step ? :D

OpenStudy (trojanpoem):

You mean the "derivative" answer ?

ganeshie8 (ganeshie8):

whichever makes sense

OpenStudy (trojanpoem):

The derivative is wrong according to my teacher. Logs maybe used. On my side , I tried it but achieved nothing. Give it a try. ::: Got some network connection problems.

ganeshie8 (ganeshie8):

\(\large n^{n!} = n\times n\times \cdots \times \text{n! times}\) \(\large n!^n = n^2\text{terms each of which are }\le n\)

ganeshie8 (ganeshie8):

since \(\large n! \gt n^2\) when \(n\ge 4\), the value of \(n^{n!}\) will be greater than \(n!^n\)

ganeshie8 (ganeshie8):

thats the first reply and it kindof makes sense right ?

OpenStudy (trojanpoem):

Maybe , but how will it help in finding the value of x ?

OpenStudy (trojanpoem):

I can be used in proving that n!^{n} is greater than n^{n!}

ganeshie8 (ganeshie8):

a > b what does that tell you ?

ganeshie8 (ganeshie8):

it says that a is NEVER equal to b right ? proving n^n! > n!^n tells us the same : there are no solutions for n >= 4

OpenStudy (trojanpoem):

You mean if we proved that a > b. It tells us that n got no possible solutions ?! or I misunderstood

ganeshie8 (ganeshie8):

Exactly ^

ganeshie8 (ganeshie8):

if you know that \(\large n^{n!} \gt n!^n\) for all \(\large n \ge 4\), doesnt that prove there are no solutions for \(\large n\ge 4\) ?

OpenStudy (trojanpoem):

My teacher said " it can be solved by 4 different methods" He even told me that the answers are ( 0, 1, 2 ) but 0 is refused.

OpenStudy (trojanpoem):

pelletTY network connection -_-

ganeshie8 (ganeshie8):

there could be more ways to approach this problem, are you trying to find atleast four different methods to sovle this ?

OpenStudy (ikram002p):

hmm i have an idea , it might not work though ..

OpenStudy (trojanpoem):

I am trying to find atleast one LAWL. What's your idea ??

OpenStudy (ikram002p):

ok x!= nx right ? thus x^nx=(x!)^x (x^n)^x=(x!)^x x^n=x! gives the only solution

OpenStudy (ikram002p):

how ever , for x>2 x^n = x! x.x...x = 1.2.3...x

OpenStudy (ikram002p):

seems like for x! >2 there is at least 2x number of primes ( can't remember how much lol ) but this should solve it :O

OpenStudy (trojanpoem):

I got a few questions, x! = nx : n stands for ?? i mean 4! = 4*3*2*1 do you mean n = 3*2*1 for ex ?

OpenStudy (ikram002p):

x!=1.2.3.4..x let 1.2.3.4...x-1 =n x!=nx ( means x divides x! )

OpenStudy (trojanpoem):

x^n = x! ok , now what will you do on this ?

OpenStudy (trojanpoem):

n = (x-1)!

OpenStudy (ikram002p):

no note this x^n means its an x terms only a number that only have x

OpenStudy (trojanpoem):

Hmm , clarify it a bit

OpenStudy (ikram002p):

@ganeshie8 help ?xD

ganeshie8 (ganeshie8):

im happy with the first solution in mse

OpenStudy (ikram002p):

ohk :O

OpenStudy (trojanpoem):

To me , it looks like the legendary mathematical problem won't be solved. All the answers are sticking to the same approach giving no forward answer :/

OpenStudy (zarkon):

\[x\approx2.4278\]

OpenStudy (trojanpoem):

(2.4278)! on calculator = Math error, am I mistaken ? -> x! <- in this case x can only be an integer number. Correct to me if I am missing something.

OpenStudy (zarkon):

\[x!=\Gamma(x+1)\]

OpenStudy (zarkon):

http://en.wikipedia.org/wiki/Gamma_function

OpenStudy (trojanpoem):

Indeed, I have no objection to your answer. But I didn't take the Gamma function at least till now. So I can't use it on exams :/. But I am still confused how can x! be 2.4 ?? When I first took the lesson the prof said "Only integers!".

OpenStudy (trojanpoem):

Can someone solve it using Logs please ?

OpenStudy (kainui):

Here we go... \[\LARGE x!^x=x^{x!}\]\[\LARGE x!=x^{(x-1)!}\] So this means x! is also just a single number raised to a power. But factorials are consecutive numbers multiplied together, so obviously the only possible choices are for when you just begin. This immediately allows 1 and 2 in because everything can be multiplied by 1, so it doesn't take away 2's "pure" number to a power status. After that, there's no way it can be satisfied.

OpenStudy (kainui):

To clarify a bit, the left side of the equation says: \[\LARGE x*(x-1)*(x-2)*...\] while the right side says \[\LARGE x*x*x*...\] So the only case it works looks like this (ok, I mean it's sort of half plugged in already, but you can see what the pattern is I'm considering here maybe more clearly by looking. \[\LARGE 1=1^{0!}\] and \[\LARGE 2*(2-1)=2^{1!}\] It's quite obvious that this and higher orders won't work I think by looking at the exponents. \[\LARGE 3*(3-1)*(3-2)=3^{2!}\]

OpenStudy (freckles):

\[x^{x!}=x^\frac{x!}{2}x^\frac{x!}{2}=(x^x)^\frac{(x-1)!}{2}(x^\frac{(x-1)!}{2})^x \\ x!^x=(x \cdot (x-1))!^x=x^{x}(x-1)!^x \\ \text{ it is obvious } \frac{(x-1)!}{2} \ge 1 \text{ for all integer } x \ge 3 \\ \text{ so we need to prove } (x-1)! \le x^\frac{(x-1)!}{2} \text{ for } x \ge 3 \\ \text{ this isn't so bad to prove by induction }\]

OpenStudy (trojanpoem):

Kainui , right, x is a single number raised to a power which does nothing (1! = 1, 0! = 1).

OpenStudy (trojanpoem):

but all this is theoretical, I wanna prove it mathematically.

OpenStudy (trojanpoem):

freckles, fine, could you try to do so ?. I failed to prove anything at least on this one :(

OpenStudy (freckles):

\[(3-1)!=2 \le 3=3^\frac{(3-1)!}{2} \\ \text{ Suppose} (x-1)! \le x^\frac{(x-1)!}{2} \text{ for some integer } x \ge 3 \\ \text{ We want to show } x! \le (x+1)^{\frac{x!}{2}} \\ x(x-1)! \le x \cdot x^\frac{(x-1)!}{2} \text{ multiply x on both side the induction hypothesis } \\ x! \le x^{\frac{(x-1)!+2}{2}} \\ (x-1)! \ge 2 \text{ for all integer } x \ge 3 \text{ so we have} \\ x! \le x^{\frac{(x-1)!+(x-1)!}{2}} =x^\frac{2(x-1)!}{2} \\ x! \le x^\frac{x(x-1)!}{2} \text{ since } x \ge 3 \\ x! \le x^\frac{x!}{2} \\ \text{ finally } x+1>x \text{ for all integer } x \ge 3 \\ x! \le (x+1)!^{\frac{x!}{2}}\] Now think everything I did seems fine.

OpenStudy (trojanpoem):

[Math Processing Error] I can't view your text. What do you mean by "seems fine" ?

OpenStudy (kainui):

"but all this is theoretical, I wanna prove it mathematically." I don't understand how my argument and reasoning is somehow not good enough? Where does your doubt come in? What more do you need to see formalized to make you feel better about it? It will only ever continue to diverge as I showed, but we can wrap it up with a lot of fancy mathematical analysis symbols but it would come down to saying exactly the same thing in only a more convoluted form.

OpenStudy (trojanpoem):

Sis bro whatever , I didn't refuse your argument , but do you think it will be wise of me to write something like that in an exam ? Of course not. Even through , All the symbols will prove what you have proved theoretically. It's obvious if x^x! = x!^x then x = x! and x , x! >= 2 but how to prove that mathematically ? that's the matter.

OpenStudy (trojanpoem):

At least , thanks for tying to help.